Calculate the de Broglie wavelength for a proton with a velocity 4.13% of the speed of light.

m
1
Expert's answer
2020-05-19T08:45:34-0400


h = 6.626 × 10-34

v = 4.13/100 * 299 792 458 = 12381428.5 m/s

m (proton) = 1.67 × 10-27 kg


λ  = 6.626 × 10-34 / (1.67 × 10-27 * 12381428.5) = 3.2 x 10-14 m